OpenStudy (anonymous):

Divide. Then simplify by taking roots, if possible. Assume all expressions under radicals represent positive numbers. sqrt400/2sqrt2

OpenStudy (anonymous):

\[\frac{\sqrt{400}}{2\sqrt{2}}\] \[\frac{20}{2\sqrt{2}}\] \[\frac{10}{\sqrt{2}}\]

OpenStudy (anonymous):

if you want to "rationalize the denominator" you can write \[5\sqrt{2}\]